# Accreditation Against Limited Adversarial Noise

### v3

@@ I Introduction
Accreditation is a type of verification that provides an efficient and scalable method for quantifying the quality of specific quantum computations, without trusting any aspect of those computations. Verification is a vital requirement for using quantum computers – and demonstrating quantum advantage – in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment (known as noise [2]) that induce erroneous operators (known as error) in a computation.
I note that accreditation is not the only method of evaluating quantum computations or quantum computers. One class of alternative protocols is randomized benchmarking [3, 4, 5, 6]. Randomized benchmarking protocols do not aim to measure how well a quantum computer can implement a specific computation (e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7]) but rather how well it performs benchmarking computations (which usually comprises a series of gates that, if implemented with no error, return a known measurement outcome with certainty) and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented and if specific data obtained from quantum computers can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and – via some machinations – decide whether to accept or reject [^1] On the grounds that noise has ruined the computation. the returned outputs. This binary measure is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as in Def. 1) which accreditation protocols provide.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework/setting of delegated computation [12] (i.e. is based on two characters Alice and Bob, each with their own opposing aims) interacting and precluding error typically [^2] There exist verification protocols that do not assume there is no error in some particular aspect of a computation, using two non-communicating quantum computers [13, 14, 15], but here I focus on single-device protocols. either in state preparation [16, 17, 18] or measurement [19, 20, 14, 21]); accreditation has instead tended to base its assumptions on the experimentally-observed physics within quantum computers. By turning from the cryptographic and trust-based assumptions of verification protocols [8] – from which it originates – to physics-based assumptions, accreditation has been able to develop assumptions and then protocols [22] – using those assumptions – to initiate a vein of research that has thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [23], and led to the first methods for quantifying the quality of the outputs of quantum analogue simulations [24, 25].
@@
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^3] Meaning in each execution of a circuit, the error is represented by identical and independent – across multiple executions – completely positive trace preserving maps (as defined in Def. 5) acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [23] to assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [12], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in – still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge (of the circuits to be executed) and abilities (to influence the execution of circuits) by adding a new, impartial, character, Robert, who actually perfroms all quantum computations. This upgrade is summarized in Table 1 and improves upon the IID CPTP error model, in Ref. [23], while retaining validity in all situations where it was valid before. The result is a protocol with less stringent assumptions, loosening the requirements on the physical computers used to implement it.
@@ II Problem Setting of This Paper
The foundation of my problem setting [^4] Which can be seen as akin to the error model in previous accreditation protocols [22, 23, 24, 25]. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [23] and depicted in Fig. 3. CPTP maps are as defined in Def. 5, which first requires several definitions.
The first of these definitions is Def. 2.
Definition 2.
Any linear operator, $\hat{A}$, acting on any Hilbert space is positive if, for any element, $|x\rangle$, in the Hilbert space:
$\langle x|\hat{A}|x\rangle\geq 0.$ (1)
Def. 2 then enables me to define a positive map, in Def. 3.
Definition 3.
A map is positive if positive operators are mapped exclusively to positive operators by it. If, $\forall N\in\mathbb{N}$, $\Phi\otimes I_{N}$ is positive (where $I_{N}$ is the identity on $N$ qubits), then $\Phi$ is completely positive.
With one half of the attributes of CPTP maps defined, I turn to the other half, and define trace-preserving maps in Def. 4.
Definition 4.
A map, $\Phi$, is trace-preserving if for any density matrix, $\rho$:
$\operatorname{Tr}\bigg{[}\Phi\big{(}\rho\big{)}\bigg{]}=\operatorname{Tr}\bigg{(}\rho\bigg{)}.$ (2)
For completeness, once completely-positive maps and trace-preserving maps have been defined, I formally define CPTP maps.
Definition 5.
A completely positive trace-preserving map (i.e. a CPTP map) is a map that is both:
$\bullet$ Completely Positive.
$\bullet$ Trace-Preserving.
@@ II.2 Problem Setting Preliminaries and Definitions
Presenting Protocol 1 first requires I define a number of concepts. These start with Def. 6, Def. 7, and Def. 8. Which, collectively, define a concept called redaction, and its usage, which is designed to model – in the problem setting – the idea that noise is independent of which single-qubit gates are applied in a circuit but can depend on all other aspects of a circuit. As noise is adversarial in this paper, to model this I need some way of showing the adversarial noise (i.e. Bob) the circuit but hiding the single-qubit gates from it/him. Redaction is how I achieve this. It is not a physical thing we actually can do but is part of the new cryptographic model I develop herein – and represents single-qubit gates experiencing gate-independent error (as the gates are hidden when the CPTP maps applying the error are chosen by Bob).
@@ II.2.2 CPTP Lists and Error-Related Definitions
When Bob wants to provide Robert with CPTP maps that implement the error in a specific execution, he gives Robert a CPTP list, as defined in Def. 9. But Bob cannot just give Robert any CPTP list he likes, there are limits imposed by Protocol 1 (and rigidly enforced by Robert) that the CPTP lists Bob provides must conform to:
@@
    The CPTP list must match (as defined in Def. 10) the circuit it will be applied in the execution of (basically meaning the CPTP map is capable of defining the error in the circuit execution).
@@
    All CPTP lists Bob provided in a single instance of Protocol 1 must be from a single Set of Probabilistically Similar CPTP Lists with parameter $\beta$ (defined in Def. 12), declared at the beginning of Protocol 1, for a $\beta\in\mathbb{R}$ known before the start of the protocol.
@@
However, a CPTP list is just a list of CPTP maps. In order to to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps and each must act on exactly the right number of qubits. In this case the CPTP list is said to fit the circuit, as in Def. 10.
@@
A key metric of error occurring in a circuit execution is defined in Def. 11, which allows for a useful quantification of the effect of the error.
@@
For any execution of a circuit afflicted by only stochastic Pauli error, the probability of error of that execution is the probability that the execution does not provide a sample from the same probability distribution that executing the circuit as intended (i.e. with no error applied) would.
I.e. it is the probability that none of the stochastic Pauli error channels in the circuit execution apply a Pauli gate. For example, if, for all density matrices, $\rho$:
@@
[…]
Having defined the probability of error in an execution of a circuit, which is a consequence of the CPTP list determining the error in that execution, I can define a formal concept, in Def. 12, wherein a set of CPTP maps that may govern error in circuit executions all have a similar probability of error.
@@
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (abbreviated as a SPSCLβ) is a set of CPTP lists, that all fit the same set of circuits (as in Def. 10), such that, for any CPTP list in the set, the probability of error in any circuit execution it determines the error for, once it is twirled to stochastic Pauli error, is within the interval:
@@
For an example of a SPSCLβ, consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 11). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as an SPSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this SPSCLβ is $0.5$).
@@ II.2.3 Relating CPTP Lists and Redaction Classes
As in Note 1, if a given CPTP list fits one circuit in a redaction class, it fits every circuit in that redaction class.
As in Def. 12 if one CPTP list in a given SPSCLβ fits a specific circuit then every CPTP list in that SPSCLβ fits that circuit.
Lemma 2 follows from combining these two facts. ∎
@@ II.3 Problem Setting Presentation
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1. Both Alice and Bob are required to participate in Protocol 1 and both must abide by all requirements of it. These requirements are enforced by another character, Robert, who actually implements the quantum computations – in Protocol 1 – according to Alice and Bob’s combined instructions. However, Protocol 1 still leaves room for Bob to attempt to trick Alice.
@@
The aims of Alice and Bob remain the same as in the standard cryptographic setting and as discussed in Sec. II.1: Alice is attempting to get Robert (via Protocol 1) to perform a quantum computation and provide her with the true results, while Bob is trying to trick Alice by adding error to Alice’s requested circuit (when Robert executes it in Protocol 1) so that Robert provides Alice with the results of an incorrect computation and Alice believes that the computation was performed without error. I.e. that the results Alice receives are the outputs of the computation she wanted to be performed when they are not (due to the error Bob adds). However, in trying to achieve these aims, Alice and Bob can only communicate via Protocol 1 with Robert. Robert’s aims are entirely neutral, he has no preference on if Alice or Bob successfully achieves their aim. He just wants to do his duty and perform Protocol 1 correctly.
The problem setting of this paper is summarized in Table 2.
@@ II.4 Physical Justification of the Problem Setting
Both of these limitations are applied (in the case of the first) or enforced (in the case of the second) by Robert.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [22, 23, 24]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [30, 31] (which has held true over time [32, 33], and can be most clearly seen in Ref [34, Fig. 5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [35] and has seen extensive use in theoretical work [3, 36, 37, 38, 39, 40, 41, 42, 43, 44, 35].
@@
    The hardware executing the similar circuits is the same in each execution and any error-inducing aspects of the hardware are unlikely to change much in the short time between executions.
@@
    Ref. [23, Figure. 5] ran the accreditation protocol developed therein many times, producing a probability of error in the trap circuit executions for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
@@
    Ref. [45, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [34, Fig. 9] examined NISQ computers and plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be – with some notable exceptions – small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
@@
    Ref. [46, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that (again, over a time-span of days) for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
@@
To my knowledge, I am the first to explicitly state this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [46, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.

### v4

@@ I Introduction
Without methods to assess the quality of quantum computers and computations, they will be untrustworthy and thus almost completely useless. There are a range of methods for establishing trust in quantum computers and computations, each approaching the problem from a different perspective. Together, these are known as quantum characterization, verification, and validation [1] (often abbreviated QCVV). QCVV can be decomposed broadly into:
  1. 1.
    Benchmarking [2, 3], which investigates and quantifies the performance of the quantum computer itself. Particularly popular is randomized benchmarking [4, 5, 6, 7].
  2. 2.
    The characterization of quantum devices [1], which seeks descriptions of the behavior of quantum devices and their capabilities.
  3. 3.
    Verification [8, 9, 10, 11], which investigates and quantifies how well specific computations [^1] E.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [12]. are performed.
All of these approaches are vital to the future development and usage of quantum computers, particularly in the NISQ era [13] – when quantum computations will be unreliable due to interactions with the environment [14] that induce erroneous operators in a computation – but will remain so into the fault-tolerant era, to quantify the quality of the outputs and monitor the performance of the error correction.
However, this paper – from here on – focuses exclusively on verification, particularly near-term applicable verification methods that quantify the effect of error rather than just detecting if error occurs. Among such verification protocols, two stand out: accreditation and mirror circuit fidelity estimation [15].
Typically, verification protocols run a quantum computation, allowing for the possibility of error, and – via some machinations – decide whether to accept or reject based on if error is detected at all. However, detecting only if error occurs is a much more limited measure of how well a specific computation is performed than either the ideal-actual variation distance (as in Def. 1), which accreditation protocols provide, or fidelity, which mirror circuit fidelity estimation provides. As such, these protocols will be more useful in the near-term, when some error will almost always occur (in large computations), and will allow for more nuanced decisions on how useful results are based on how severe the error is.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation and mirror circuit fidelity estimation. Verification has typically used the cryptographic framework/setting of delegated computation [16] (i.e. is based on two characters Alice and Bob, each with their own opposing aims) interacting and precluding error typically [^2] There exist verification protocols that do not assume there is no error in some particular aspect of a computation, using two non-communicating quantum computers [17, 18, 19], but herein I focus on single-device protocols. either in state preparation [20, 21, 22] or measurement [23, 18, 24]); accreditation and mirror circuit fidelity estimation have instead tended to base their assumptions on the experimentally-observed physics within quantum computers.
By turning from the cryptographic and trust-based assumptions of verification protocols [8] – from which it originates – to physics-based assumptions, accreditation has been able to develop assumptions and then protocols [25] – using those assumptions – to initiate a vein of research that has thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [26], and led to the first methods for quantifying the quality of the outputs of quantum analogue simulations [27, 28].
@@
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^3] Meaning in each execution of a circuit, the error is represented by identical and independent – across multiple executions – completely positive trace preserving maps acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [26] to assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [16], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in – still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge (of the circuits to be executed) and abilities (to influence the execution of circuits) by adding a new, impartial, character, Robert, who actually perfroms all quantum computations. This upgrade is summarized in Table 1 and improves upon the IID CPTP error model, in Ref. [26], while retaining validity in all situations where it was valid before. The result is a protocol with less stringent assumptions, loosening the requirements on the physical computers used to implement it.
@@ II Problem Setting of This Paper
The foundation of my problem setting [^4] Which can be seen as akin to the error model in previous accreditation protocols [25, 26, 27, 28]. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [26] and depicted in Fig. 3.
@@ II.2 Problem Setting Preliminaries and Definitions
Presenting Protocol 1 first requires I define a number of concepts. These start with Def. 2, Def. 3, and Def. 4. Which, collectively, define a concept called redaction, and its usage, which is designed to model – in the problem setting used herein – the idea that noise is independent of which single-qubit gates are applied in a circuit but can depend on all other aspects of a circuit. As noise is adversarial in this paper, to model this I need some way of showing the adversarial noise (i.e. Bob) the circuit but hiding the single-qubit gates from it/him. Redaction is how I achieve this. It is not a physical thing we actually can do but is part of the new cryptographic model I develop herein – and represents single-qubit gates experiencing gate-independent error (as the gates are hidden when the CPTP maps applying the error are chosen by Bob).
@@ II.2.2 CPTP Lists and Error-Related Definitions
When Bob wants to provide Robert with CPTP maps that implement the error in a specific execution, he gives Robert a CPTP list, as defined in Def. 5. But Bob cannot just give Robert any CPTP list he likes, there are limits imposed by Protocol 1 (and rigidly enforced by Robert) that the CPTP lists Bob provides must conform to:
@@
    The CPTP list must match (as defined in Def. 6) the circuit it will be applied in the execution of (basically meaning the CPTP map is capable of defining the error in the circuit execution).
@@
    All CPTP lists Bob provided in a single instance of Protocol 1 must be from a single Set of Probabilistically Similar CPTP Lists with parameter $\beta$ (defined in Def. 8), declared at the beginning of Protocol 1, for a $\beta\in\mathbb{R}$ known before the start of the protocol.
@@
However, a CPTP list is just a list of CPTP maps. In order to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps and each must act on exactly the right number of qubits. In this case the CPTP list is said to fit the circuit, as in Def. 6.
@@
A key metric of error occurring in a circuit execution is defined in Def. 7, which allows for a useful quantification of the effect of the error.
@@
For any execution of a circuit afflicted by only stochastic Pauli error, the probability of error of that execution is the probability that any Pauli gate is erroneously applied at any point in the execution by an error channel. I.e. it is the probability that none of the stochastic Pauli error channels in the circuit execution apply a Pauli gate.
For example, if, for all density matrices, $\rho$:
@@
then the probability of error in the execution of a circuit where these channels model the error (e.g. Fig. 3) is the probability that any of $\xi_{1}$, $\xi_{2}$, or $\xi_{3}$ apply error to the circuit execution ($1-\frac{1}{2}\frac{1-0.01}{2}\frac{1+0.01}{2}\approx 0.875$).
Having defined the probability of error in an execution of a circuit, which is a consequence of the CPTP list determining the error in that execution, I can define a formal concept, in Def. 8, wherein a set of CPTP maps that may govern error in circuit executions all have a similar probability of error.
@@
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (abbreviated as a SPSCLβ) is a set of CPTP lists, that all fit the same set of circuits (as in Def. 6), such that, for any CPTP list in the set, the probability of error in any circuit execution it determines the error for, once it is twirled to stochastic Pauli error, is within the interval:
@@
For an example of a SPSCLβ, consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 7). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as an SPSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this SPSCLβ is $0.5$).
@@ II.2.3 Relating CPTP Lists and Redaction Classes
As in Note 1, if a given CPTP list fits one circuit in a redaction class, it fits every circuit in that redaction class. As in Def. 8, if one CPTP list in a given SPSCLβ fits a specific circuit then every CPTP list in that SPSCLβ fits that circuit. Lemma 2 follows from combining these two facts. ∎
@@ II.3 Problem Setting Presentation
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1. Both Alice and Bob are required to participate in Protocol 1, when initiated, and both must abide by all requirements of it at a. These requirements are enforced by another character, Robert, who actually implements the quantum computations – in Protocol 1 – according to Alice and Bob’s combined instructions. However, Protocol 1 still leaves room for Bob to attempt to trick Alice.
@@
The aims of Alice and Bob remain the same as in the standard cryptographic setting and as discussed in Sec. II.1: Alice is attempting to get Robert (via Protocol 1) to perform a quantum computation and provide her with the true results, while Bob is trying to trick Alice by adding error to Alice’s requested circuit (when Robert executes it in Protocol 1) so that Robert provides Alice with the results of an incorrect computation and Alice believes that the computation was performed without error. I.e. that the results Alice receives are the outputs of the computation she wanted to be performed when they are not (due to the error Bob adds). However, while trying to achieve these aims, Alice and Bob can only communicate via Protocol 1 – with Robert. Robert’s aims are entirely neutral, he has no preference on if Alice or Bob successfully achieves their aim. He just wants to do his duty and perform Protocol 1 correctly. The problem setting of this paper is summarized in Table 2.
@@ II.4 Physical Justification of the Problem Setting
Both of these limitations are applied (in the case of the first) or enforced (in the case of the second), in Protocol 1, by Robert.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [25, 26, 27]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [33, 34] (which has held true over time [35, 36], and can be most clearly seen in Ref [37, Fig. 5]): the error is so small that the error in different single-qubit gates cannot differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [38] and has seen extensive use in theoretical work [4, 39, 40, 41, 42, 43, 44, 45, 46, 47, 38].
@@
    The hardware executing the similar circuits is the same in each execution and any error-inducing aspects of the hardware, or surrounding environment, are unlikely to change much in the short time between executions.
@@
    Ref. [26, Figure. 5] ran the accreditation protocol developed therein many times, producing a probability of error in the trap circuit executions for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
@@
    Ref. [48, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [37, Fig. 9] examined NISQ computers and plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be – with some notable exceptions – small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
@@
    Ref. [49, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that (again, over a time-span of days) for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
@@
To my knowledge, I am the first to explicitly state this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [4, 5, 6, 7] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [49, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.